OpenStudy (here_to_help15):

"Include three copied proofs with their diagrams (yes, that's right... copy and paste!) of the Pythagorean Theorem, with sources cited (use the Internet, your textbook, other math textbooks, etc). They must be PROOFS! This is different than examples using the Pythagorean Theorem. You do NOT need to explain the proofs in your own words."

OpenStudy (here_to_help15):

My teachers Instructions : )

OpenStudy (acxbox22):

u kno u can use google rite?

OpenStudy (here_to_help15):

Im having trouble locating this on google is why

OpenStudy (here_to_help15):

Can you list me sites please?

OpenStudy (here_to_help15):

@acxbox22

OpenStudy (here_to_help15):

ΔABF = ΔAEC by SAS. This is because, AE = AB, AF = AC, and ∠BAF = ∠BAC + ∠CAF = ∠CAB + ∠BAE = ∠CAE. This is one proof corrects?

OpenStudy (acxbox22):

there are many proofs, look through the websites because they all have different proofs which you cant copy/paste but can type up yourself

OpenStudy (here_to_help15):

What i circled that is a proof correct?

OpenStudy (acxbox22):

no that is part of the proof the proofs are pretty long but copy the text and copy the pictures and try to understand them

OpenStudy (here_to_help15):

wait dont go one sec

OpenStudy (acxbox22):

ik thanks for the testi :)

OpenStudy (here_to_help15):

Thats the whole proof correct?

OpenStudy (acxbox22):

copy all the way to the part it says proof #2 thats how you know the proof has ended

OpenStudy (here_to_help15):

lol can you do it ? I am not understanding you ;) here is the link i got this from http://www.cut-the-knot.org/pythagoras/

OpenStudy (here_to_help15):

And no problem :) my pleasure :D

OpenStudy (here_to_help15):

Bhaskara's proof is also a dissection proof. It is similar to the proof provided by Pythagoras. Bhaskara was born in India. He was one of the most important Hindu mathematicians of the second century AD. He used the following diagrams in proving the Pythagorean Theorem. In the above diagrams, the blue triangles are all congruent and the yellow squares are congruent. First we need to find the area of the big square two different ways. First let's find the area using the area formula for a square. Thus, A=c^2. Now, lets find the area by finding the area of each of the components and then sum the areas. Area of the blue triangles = 4(1/2)ab Area of the yellow square = (b-a)^2 Area of the big square = 4(1/2)ab + (b-a)^2 = 2ab + b^2 - 2ab + a^2 = b^2 + a^2 Since, the square has the same area no matter how you find it A = c^2 = a^2 + b^2, concluding the proof.

OpenStudy (here_to_help15):

Hehe thanks i already got it : )

OpenStudy (here_to_help15):

@acxbox22 ok one more question and im done : )